The following are the transactions relating to the formation of Cardinal Mowing Services, Inc., and its first month of operations. a. The firm was organized and the initial stockholders invested cash of $840 b. The company borrowed $1,260 from a relative of one of the initial stockholders, a short-term note was signed c. Two zero-turn lawn mowers costing $672 each and a professional trimmer costing $182 were purchased for cash. The original list price of each mower was $854, but a discount was received because the seller was having a sale d. Gasoline, oil, and several packages of trash bags were purchased for cash of $126. e. Advertising flyers announcing the formation of the business and a newspaper ad were purchased. The cost of these items, $238, will be paid in 30 days f. During the first two weeks of operations, 47 lawns were mowed. The total revenue for this work was $987, $651 was collected in cash, and the balance will be received within 30 days. g. Employees were paid $588 for their work during the first two weeks h. Additional gasoline, oil, and trash bags costing $154 were purchased for cash i. In the last two weeks of the first month, revenues totaled $1,288, of which $525 was collected. j. Employee wages for the last two weeks totaled $714; these will be paid during the first week of the next month k. It was determined that at the end of the month the cost of the gasoline, oil, and trash bags still on hand was $42 L. Customers paid a total of $210 due from mowing services provided during the first two weeks. The revenue for these services was recognized in transaction f

a. Record each transaction in the appropriate columns. (If an transaction/Adjustment are not affecting the balance sheet category or income statement, leave the cells blank. Enter decreases to account balances as a negative.) STOCKHOLDERS EQUITY ITIE ASSETS Retained Accounts Paid in Notes Accounts Transaction ble CapitalEarningsRevenueExpenses Cash +Receivable Supplies+EquipmentPayable 840 1.260 1,260+ (1.526) 1.526 C. (126) | + 126ク+ 238ク 987クー 651 1,638 588ク (588) + (154) 154> + 1,288 이- 7140 238 525+ 763 714 238) k. 1,778 (210) | + 2,191 + 2.275 210+ 1,092 + 1,526 1,260+ 952 + 840+ 42 + no response was expected in cel or a formula-based calculat on。incorrectJ no points deducted. Rad taxt indicates

c. After completing parts a through I, prepare an income statement for Cardinal Mowing Services, Inc., for the month presented and a balance sheet at the end of the month. (Enter decreases to account balances as a negative.) CARDINAL MOWING SERVICES, INC. Income Statement |$ Service revenue Advertising expense Supplies expense Wages expense 1,687 3 238) (238) (714) Net income S 497

CARDINAL MOWING SERVICES, INC. Balance Sheet Assets: Cash Accounts receivable Supplies Equipment 1,092 2,191 42 1,526 Total current assets 4,851 Total assets 4,851 Liabilities Notes payable Accounts payable S 1,260 952 otal liabilities $ 2,212 Stockholders Equity Paid-in-capital 840 840 Total stockholders equity Total liabilities and stockholders $ 3,052
Answer #1

A. Recording transactions:-

Assets Liabilities Stockholders Equity
Transaction Cash + Accounts receivable + Supplies + Equipment = Notes payable + Accounts Payable + Paid - in - capital + Retained earnings + Revenue - Expenses
a 840 + + + = + + 840 + + -
b 1,260 + + + = 1,260 + + + + -
c (1,526) + + + 1,526 = + + + + -
d (126) + + 126 + = + + + + -
e + + + = + 238 + + + - 238
f 651 + 336 + + = + + + + 987 -
g (588) + + + = + + + + - 588
h (154) + + 154 + = + + + + -
i 525 + 763 + + = + + + + 1,288 -
j + + + = + 714 + + + - 714
k + + (238) + = + + + + - 238
l 210 + (210) + + = + + + + -
Balance 1,092 + 889 + 42 + 1,526 = 1,260 + 952 + 840 + + 2,275 - 1,778

CARDINAL MOWING SERVICES INC.

Income Statement

Items Amount
Service Revenue 2,275
Less:- Expenses:-
Advertising expenses (238)
Supplies expenses (238)
Wages Expenses (588+714=1,302) (1,302)
Net Income 497

CARDINAL MOWING SERVICES INC.

Balance Sheet

Items Amount
Assets:-
Cash 1,092
Accounts Receivable 889
Supplies 42
Equipment 1,526
Total Assets 3,529
​​​​​​Liabilities:-
Notes Payable 1,260
Accounts Payable 952
Total Liabilities 2,212
Stockholders Equity:-
Paid - in - capital 840
Retained Earnings 497
Total Stockholders Equity 1,337
Total Liabilities & Stockholders Equity 3,549